38 year old man has an 11 year old daughter. Someday he will be twice as old as his daughter. How old will she be when this occurs?
Found 3 solutions by stanbon, rwm, Alan3354:
Answer by stanbon(75887)

38 year old man has an 11 year old daughter. Someday he will be twice as old as his daughter. How old will she be when this occurs?
---------------
Now:
father = 38
daughter = 11
---------------------
Someday:
father = 38+x
daughter = 11+x
---------------------------
Equation:
(38+x) = 2(11+x)
38 +x = 22 + 2x
15 = x
----
Daughter's ang then: 11+x = 11+15 = 26
==================================
Cheers,
Stan H.

Answer by rwm(914)

38+x=2*(11+x)
38+x=22+2x
16=x
she will be 11+16=27
He will be 54=38+16

Answer by Alan3354(69443)

The difference is 27 yrs, and will always be 27 yrs.
--> when she's 27.